How do you solve an eqation like this? 4x-5(3x+10)=126
Answer by nerdybill(7384)

4x-5(3x+10)=126
First, distribute the -5 to terms inside the parenthesis:
4x-15x-50 = 126
combine like-terms:
-11x-50 = 126
add 50 to both sides:
-11x = 176
finally, divide both sides by 11
x = -176/11
x = -16